40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 15 01:14:10 2015, rana, Summary, LSC, need to check LSC Whitening switch logic ... again 
    Reply  Wed Sep 16 15:04:33 2015, ericq, Summary, LSC, DC PD Whitening Board Fixed 
Message ID: 11606     Entry time: Wed Sep 16 15:04:33 2015     In reply to: 11597
Author: ericq 
Type: Summary 
Category: LSC 
Subject: DC PD Whitening Board Fixed 
Quote:

Tonight we noticed that the REFL_DC signal has gone bipolar, even though the whitening gain is 0 dB and the whitening filter is requested to be OFF.

Fixed! I noticed that whitening gain changes weren't having any effect on CM_SLOW. I then checked REFL_DC, where this also seemed to be the case. Since the gain is controlled via VME machine, and whitening filter switching is controlled via RCG, I figured there must be something wrong with the board. I checked all of the DC PD signals, which share a whitening filter board, and they all had the same symptoms. 

I went and peeked at the board, and it turns out the backplane cable had fallen off. frown

I plugged it in, things look ok. 

ELOG V3.1.3-